Given that f(x)=8x^2+7 and g(x)=4x+4 are functions from R to R, find
(a) f∘g.
(b) g∘f.
1
Expert's answer
2020-10-28T16:42:39-0400

"f(x)=8x^2+7" "g(x)=4x+4"


a) "f\circ g=f(g(x))=8(4x+4)^2+7=128x^2+256x+135"


b) "g \circ f=g(f(x))=4(8x^2+7)+4=32x^2+32"